# Absolute continuity of non-Gaussian and Gaussian multiplicative chaos measures

### v1

@@ 2.1. The discrete hierarchical model
As an immediate consequence of Lemma 2.2 and the fact that both $\operatorname{Var}[\widetilde{S}_{n,\widetilde{a}}(t)]$, $\operatorname{Var}[S_{2^{n}-1,a}(t)]$ may be expressed as $n\log(2)+o_{n}(1)$ we get the following estimate on the partition functions: for each $t\in[0,1]$ and $\gamma\in(0,\sqrt{2})$, there exist $\zeta(\gamma,t)$ and $\widetilde{\zeta}(\gamma,t)$ such that, for all $n\in\mathbb{N}$,
$Z_{2^{n}-1,\gamma,a}(t)=\exp\Big{(}\frac{\gamma^{2}}{2}\log(2)n+\zeta(\gamma,t)+o_{n}(1)\Big{)}\\ \text{and}\quad\widetilde{Z}_{n,\gamma,\widetilde{a}}(t)=\exp\Big{(}\frac{\gamma^{2}}{2}\log(2)n+\widetilde{\zeta}(\gamma,t)+o_{n}(1)\Big{)}\,.$ (2.9)
@@
Proposition 2.3 below justifies the introduction of the hierarchical chaos measure. We prove it via a chaining argument.
@@
Let $\gamma\in(0,\sqrt{2})$, fix $(a_{k}^{(i)})_{k\in\mathbb{N}}$ i.i.d. with (1.2) and let $(\widetilde{a}_{n})_{n\in\mathbb{N}}$ be defined as in (2.3). Then, almost surely
$\widetilde{\mu}_{\gamma,\widetilde{a}}\ll\mu_{\gamma,a}\ll\widetilde{\mu}_{\gamma,\widetilde{a}}$
@@
By definition (1.3), (2.8) we have
$\widetilde{\mu}_{\gamma,\widetilde{a}}[A]$ $=\lim\limits_{n\to\infty}\int_{A}\widetilde{Z}_{n,\gamma,\widetilde{a}}(t)^{-1}e^{\gamma\widetilde{S}_{n,a}(v_{n}(t))}\mathrm{d}t=\lim\limits_{n\to\infty}\int_{A}R_{n,\gamma}(t)Z_{2^{n}-1,\gamma}(t)^{-1}e^{\gamma S_{2^{n}-1}(t)}\;\mathrm{d}t\,,\text{ and }$
$\mu_{\gamma,a}[A]$ $=\lim\limits_{n\to\infty}\int_{A}Z_{2^{n}-1,\gamma}(t)^{-1}e^{\gamma S_{2^{n}-1}(t)}\mathrm{d}t,$
so that to prove Proposition 2.3 it suffices to show that $\sup_{n\in\mathbb{N}}\sup_{t\in[0,1]}|\log(R_{n,\gamma}(t))|$ is almost surely finite. From (2.9), we know that $\widetilde{Z}_{n,\gamma,\widetilde{a}}(t)^{-1}Z_{2^{n}-1,\gamma,a}(t)$ converges. Thus, it suffices to show that, almost surely,
@@
Recalling the map $\pi_{n}$ from Section 2.1, along with (2.5) and (2.3), we can write
$|\widetilde{S}_{n,a}(t)-S_{2^{n}-1}(t)|=\Bigg{|}\sum_{m=1}^{n}\sum_{k=2^{m}-1}^{2^{m}-1}X_{k}(t)-X_{k}(\pi_{m}(t))\Bigg{|}\,.$
In light of Borel-Cantelli, to show (2.10), we only need to show the existence of $C_{0}>0$ such that
@@
Observe Eq. 2.7 implies $Y_{m}(t_{m}(j))=0$ for all $j\in\llbracket 1,T_{m}\rrbracket$, and so the chaining argument proceeds by considering a dyadic mesh of points within each interval $I_{m}(j)$, enumerated as follows. For $\ell,m\in\mathbb{N}$, define
@@
so that $|\widetilde{\mathcal{P}}_{\ell,m}|=T_{m}2^{\ell}$. Furthermore, let $\rho_{\ell,m}(t)$ denote the unique element $r\in\widetilde{\mathcal{P}}_{\ell,m}$ minimizing distance to $t$ and satisfying $\pi_{m}(t)=\pi_{m}(r)$ (so that $t$ and $\rho_{\ell,m}(t)$ lie in the same interval $I_{m}(j)$); in particular, $\rho_{0,m}(t)=\pi_{m}(t)$ for each $t\in[0,1]$. Since $Y_{m}(t)$ is continuous on $[0,1]\setminus\mathcal{P}_{m}$ and equal to 00 on $\mathcal{P}_{m}$, we have the following for all $t\in[0,1]$:
@@
We now record the key estimate. Consider $\ell,m\in\mathbb{N}$ and $s,t\in\widetilde{\mathcal{P}}_{\ell,m}$ that are consecutive, so that $|s-t|=2^{-\ell}|I_{m}(j)|$ for some $j\leq T_{m}$ (and therefore $\pi_{m}(t)=\pi_{m}(s)$). Then, using the Lipschitz property of cosine/sine as well as Eq. 2.1, we find
$\mathbb{P}\bigg{(}|Y_{m}(t)-Y_{m}(s)|>2\sqrt{2}\pi m^{-2}\ell^{-2}\bigg{)}=\mathbb{P}\bigg{(}\Big{|}\sum_{k=2^{m-1}}^{2^{m}-1}X_{k}(t)-X_{k}(s)\Big{|}>2\sqrt{2}\pi m^{-2}\ell^{-2}\bigg{)}\leq Ce^{-c\frac{f(m)2^{\ell}}{m^{2}\ell^{2}}},$
where we have used the exponentially-decaying right and left tails of the sum given to us by [petrov, Chapter III, Theorem 16] (in particular, Eq. (4.3) and (4.6) there). [^2] Note that [petrov, Chapter III, Lemma 5] allows us to apply Theorem 16 of that chapter, as condition (4.1) is satisfied. For each $m\in\mathbb{N}$, a union bound over $\ell\in\mathbb{N}$ and all possible pairs of $(\rho_{\ell-1,m}(t),\rho_{\ell,m}(t))$ then yields
$\mathbb{P}\Big{(}\exists_{\ell\in\mathbb{N}}\exists_{t\in[0,1]}:|Y_{m}(\rho_{\ell,m}(t))-Y_{m}(\rho_{\ell-1,m}(t))|>2\sqrt{2}\pi m^{-2}\ell^{-2}\Big{)}\leq C\sum_{\ell\in\mathbb{N}}T_{m}2^{\ell+1}e^{-c\frac{f(m)2^{\ell}}{m^{2}\ell^{2}}}.$
Combining the last display with (2.13) yields
@@
which, recalling the definition of $Y_{m}(t)$ in (2.12) as well as the bound on $T_{m}$ from (2.1), implies (2.11). This concludes the proof of Proposition 2.3. ∎
@@ 2.2. Coupling with Gaussian log-correlated fields
In addition to Proposition 2.3, the other key input for the proof of 1.1 is Proposition 2.7, which gives mutual absolute continuity of $\widetilde{\mu}_{\gamma,\widetilde{a}}$ with a chaos measure associated to a discrete hierarchical Gaussian model constructed on the same probability space. Before stating the proposition, we construct a coupling between $(\widetilde{a}_{n}^{(i)})_{i\in\{1,2\},n\in\mathbb{N}}$ and a Gaussian random vector $(\widetilde{g}_{n}^{(i)})_{i\in\{1,2\},n\in\mathbb{N}}$, where $\widetilde{g}_{n}^{(i)}:=(\widetilde{g}_{n}^{(i)}(v))_{v\in\mathcal{N}_{n}}$, such that the two vectors are close on coordinates that are “thick” (Lemma 2.6), and
@@
Note (2.14) implies “generational independence”: $\widetilde{g}_{n}^{(i)}$ is independent of $\widetilde{g}_{m}^{(i)}$ for $n\neq m$. In fact, we construct i.i.d. standard Gaussians $g_{k}^{(i)}$ such that the $\widetilde{g}_{n}^{(i)}$ are given by taking $a_{k}^{(i)}:=g_{k}^{(i)}$ in (2.3). These will be the $g_{k}^{(i)}$ appearing in the statement of 1.1. The construction is completed in (2.24). The main tool here is a variant of the Yurinskii coupling from [BELLONI20194], reproduced below.
Lemma 2.4 ([BELLONI20194, Appendix I, Lemma 38]).
@@
is finite, where each $\mathfrak{g}_{k}$ denotes an independent random vector in $\mathbb{R}^{K}$ with $\mathfrak{g}_{k}\sim\mathcal{N}(0,\operatorname{Cov}(\xi_{i}))$. Define $S:=\sum_{k=1}^{n}\xi_{k}$ and let $Z\sim\mathcal{N}(0,I_{K})$. Then, for each $\delta>0$, there exists a $\mathcal{N}(0,\operatorname{Cov}(S))$-distributed random vector $T$ such that
@@
Fix $n\in\mathbb{N}$ and $\mathcal{K}\subseteq\mathcal{N}_{n}$. There exists a $\mathcal{N}(0,\operatorname{Cov}[(\widetilde{a}_{n}^{(1)}(v),\widetilde{a}_{n}^{(2)}(v))_{v\in\mathcal{K}}])$-distributed random vector $\mathfrak{g}_{\mathcal{K},n}$ such that
@@
$\mathbb{P}\Big{(}\big{\|}\mathfrak{g}_{\mathcal{K},n}-(\widetilde{a}_{n}^{(1)}(v),\widetilde{a}_{n}^{(2)}(v))_{v\in\mathcal{K}}\big{\|}_{\infty}>n^{-2}\Big{)}\leq\min_{t\geq 0}\Big{\{}2\mathbb{P}\Big{(}\max_{j\leq K}|Z_{j}|>t\Big{)}+Cn^{6}\beta t^{2}\Big{\}},$ (2.17)
with $(Z_{j})_{j\leq K}$ i.i.d., $\mathcal{N}(0,1)$-distributed and $\beta$ as in (2.15). To control the $\xi_{k}$ part of $\beta$, we use that $\|\xi_{k}\|_{2}^{2}\leq K\|\xi_{k}\|_{\infty}^{2}$ and $\|\xi_{k}\|_{\infty}\leq N^{-1/2}\max\{|a_{l_{n}(k)}^{(1)}|,|a_{l_{n}(k)}^{(2)}|\}$ to get
$\sum_{k=1}^{N}\mathbb{E}[\|\xi_{k}\|_{2}^{2}\,\|\xi_{k}\|_{\infty}]\leq N^{-1/2}K.$ (2.18)
The $\mathfrak{g}_{k}$ part of $\beta$ is bounded as $CN^{-1/2}K\log(K)^{1/2}$ by [new-yurinskii, Lemma B.3], so that
@@
Furthermore, using Borel-TIS and $\mathbb{E}\left[\max_{j\leq K}|Z_{j}|\right]\leq C\log(K)^{1/2}$, we find for $t\geq C\log(K)^{1/2}$
$\mathbb{P}\Big{(}\max_{j\leq 2K}|Z_{j}|>t\Big{)}\leq Ce^{-\frac{(t-C\log(K)^{1/2})^{2}}{2}}.$
Plugging this into (2.17) bounds the left-hand side of (2.17) by
$C\min_{t\geq C\log(K)^{1/2}}\Big{\{}e^{-\frac{(t+C\log(K)^{1/2})^{2}}{2}}+n^{6}t^{2}N^{-\frac{1}{2}}K\log(K)^{\frac{1}{2}}\Big{\}}.$
Taking $t=C\log(K)^{1/2}+n$ while keeping in mind $N^{-1/2}=2^{-n/2}$ and $K=2|\mathcal{K}|$ yields (2.16). ∎
We now use Lemma 2.5 to recursively construct the hierarchical Gaussians $(\widetilde{g}_{n}^{(i)})_{i\in\{1,2\},n\in\mathbb{N}}$. As a first step we set $(\widetilde{g}_{1}^{(i)})_{i\in\{1,2\}}:=\mathfrak{g}_{\mathcal{N}_{1},1}$ from Lemma 2.5.
Assume we have constructed $(\widetilde{g}_{k}^{(i)})_{i\in\{1,2\},k\leq n}$. We construct the $(n+1)^{\mathrm{st}}$ generation as follows. First, we only apply the coupling on points satisfying some notion of thickness “up to level $n$”, thereby reducing the quantity of random variables we need to couple. Towards this, fix $\delta>0$ satisfying $\gamma-\delta>1$, and define the following sets of time-$n$ thick points:
$\mathcal{T}_{n,\gamma,\widetilde{a}}:=\left\{v\in\mathcal{N}_{n}:\widetilde{S}_{n,\widetilde{a}}(v)\geq\log(2)(\gamma-\delta)n\right\}$ $\,,\qquad\mathcal{T}_{n,\gamma,\widetilde{g}}:=\left\{v\in\mathcal{N}_{n}:\widetilde{S}_{n,\widetilde{g}}(v)\geq\log(2)(\gamma-\delta)n\right\}\,,$
@@
$\mathbb{P}\Big{(}\sup_{v\in\mathcal{N}_{n+1}(\mathcal{T}_{n,\gamma}),i\in\{1,2\}}\Big{|}\widetilde{g}_{n+1}^{(i)}(v)-\widetilde{a}_{n+1}^{(i)}(v)\Big{|}>n^{-2}\Big{)}\leq C2^{-\frac{(\gamma-\delta)^{2}-1}{5}n}\,.$ (2.20)
Proof.
Observe that, by construction, $\mathcal{N}_{n+1}(\mathcal{T}_{n,\gamma})$ is independent of $(\widetilde{a}_{n+1}^{(i)})_{i\in\{1,2\}}$. As a consequence, we may apply Lemma 2.5 to construct a random vector $(\widetilde{g}_{n+1}^{(1)}(v),\widetilde{g}_{n+1}^{(2)}(v))_{v\in\mathcal{N}_{n+1}(\mathcal{T}_{n,\gamma})}$ with $\operatorname{Cov}[(\widetilde{g}_{n+1}^{(1)}(v),\widetilde{g}_{n+1}^{(2)}(v))_{v\in\mathcal{N}_{n+1}(\mathcal{T}_{n,\gamma})}]=\operatorname{Cov}[(\widetilde{a}_{n+1}^{(1)}(v),\widetilde{a}_{n+1}^{(2)}(v))_{k\in\mathcal{N}_{n+1}(\mathcal{T}_{n,\gamma})}]$ and
@@
[…]
@@
We combine (2.21) and (2.22) to get (2.20). ∎
Having constructed the Gaussian vector $(\widetilde{g}_{n}^{(i)})_{v\in\mathcal{N}_{n}(\mathcal{T}_{n-1,\gamma})}$ for each $n\in\mathbb{N}$ and $i\in\{1,2\}$, we now extend this vector to $(\widetilde{g}_{n}^{(i)})_{v\in\mathcal{N}_{n}}$ in such a way that Eq. 2.14 is satisfied. Choose an arbitrary set $\overline{\mathcal{N}_{n}}$ such that $|\overline{\mathcal{N}_{n}}|=2^{n-1}$ and [^3] By a first moment bound on the right tail of $|\mathcal{N}_{n}(\mathcal{T}_{n-1,\gamma})|$, this is possible for all but finitely many $n$.
@@
Extend $(\widetilde{g}_{n}^{(i)}(v))_{v\in\mathcal{N}_{n}(\mathcal{T}_{n-1,\gamma})}$ to a Gaussian vector $(\widetilde{g}_{n}^{(i)}(v))_{v\in\overline{\mathcal{N}_{n}}}$ such that
@@
We see from Eq. 2.3 that (2.14) is satisfied. We also note that the $g_{k}^{(i)}$ constructed in Eq. 2.23 will be the ones appearing in the statement of 1.1.
As mentioned at the start of the subsection, the relationship between the $g_{k}^{(i)}$ and the $\widetilde{g}_{n}^{(i)}$ is the same as that of the $a_{k}^{(i)}$ and the $\widetilde{a}_{n}^{(i)}$, given by Eq. 2.3. As such, our construction fits in the framework of Section 2.1. By taking $a_{k}^{(i)}=g_{k}^{(i)}$ in Eq. 2.3, we may define the field $\widetilde{S}_{k,g}(v)$ as in Eq. 2.6, as well as the limiting chaos measure $\widetilde{\mu}_{\gamma,\widetilde{g}}$ as in Eq. 2.8.

### v2

@@ 2.1. The discrete hierarchical model
As an immediate consequence of Lemma 2.2 and the fact that both $\operatorname{Var}[\widetilde{S}_{n,\widetilde{a}}(t)]$, $\operatorname{Var}[S_{2^{n}-1,a}(t)]$ may be expressed as $n\log(2)+o_{n}(1)$ we get the following estimate on the sequences of normalizing functions in Eqs.˜1.3 and 2.8: for each $t\in[0,1]$ and $\gamma\in(0,\sqrt{2})$, there exist $\zeta(\gamma,t)$ and $\widetilde{\zeta}(\gamma,t)$ such that, for all $n\in\mathbb{N}$,
$Z_{2^{n}-1,\gamma,a}(t)=\exp\Big(\frac{\gamma^{2}}{2}\log(2)n+\zeta(\gamma,t)+o_{n}(1)\Big)\\ \text{and}\quad\widetilde{Z}_{n,\gamma,\widetilde{a}}(t)=\exp\Big(\frac{\gamma^{2}}{2}\log(2)n+\widetilde{\zeta}(\gamma,t)+o_{n}(1)\Big)\,$, (2.9)
where the $o_{n}(1)$ terms have no $t$ dependence since the $O_{n}$ terms in Lemma 2.2 do not.
@@
Proposition˜2.3 below justifies the introduction of the hierarchical chaos measure. We prove the proposition via a chaining argument.
@@
Let $\gamma\in(0,\sqrt{2})$, fix $(a_{k}^{(i)})_{k\in\mathbb{N}}$ i.i.d. with Eq.˜1.2 and let $(\widetilde{a}_{n})_{n\in\mathbb{N}}$ be defined as in (2.3). Then, almost surely,
$\widetilde{\mu}_{\gamma,\widetilde{a}}\ll\mu_{\gamma,a}\ll\widetilde{\mu}_{\gamma,\widetilde{a}}\,$.
@@
By the definitions in Eq.˜1.3, Eq.˜2.8 we have that for all $A\subseteq[0,1]$ Borel, $\mathbb{P}$-a.s.,
$\mu_{\gamma,a}[A]$ $=\lim\limits_{n\to\infty}\int_{A}Z_{2^{n}-1,\gamma}(t)^{-1}e^{\gamma S_{2^{n}-1,a}(t)}\mathrm{d}t\,,\quad\text{and}$
$\widetilde{\mu}_{\gamma,\widetilde{a}}[A]$ $=\lim\limits_{n\to\infty}\int_{A}\widetilde{Z}_{n,\gamma,\widetilde{a}}(t)^{-1}e^{\gamma\widetilde{S}_{n,a}(v_{n}(t))}\mathrm{d}t=\lim\limits_{n\to\infty}\int_{A}R_{n,\gamma}(t)Z_{2^{n}-1,\gamma}(t)^{-1}e^{\gamma S_{2^{n}-1,a}(t)}\;\mathrm{d}t,$
so that to prove Proposition 2.3 it suffices to show that $\sup_{n\in\mathbb{N}}\sup_{t\in[0,1]}|\log(R_{n,\gamma}(t))|$ is almost surely finite. From Eq.˜2.9, we know that $\widetilde{Z}_{n,\gamma,\widetilde{a}}(t)^{-1}Z_{2^{n}-1,\gamma,a}(t)$ converges uniformly in $t$. Thus, it suffices to show that, almost surely,
@@
Recalling the map $\pi_{n}$ from Section˜2.1, along with (2.3) and Eq.˜2.5, we can write
$|\widetilde{S}_{n,a}(t)-S_{2^{n}-1,a}(t)|=\Bigg|\sum_{m=1}^{n}\sum_{k=2^{m}-1}^{2^{m}-1}X_{k}(t)-X_{k}(\pi_{m}(t))\Bigg|\,.$
In light of Borel-Cantelli, to show Eq.˜2.10, it suffices to show the existence of $C_{0}>0$ such that
@@
Observe Eq.˜2.7 implies $Y_{m}(t_{m}(j))=0$ for all $j\in\llbracket 1,T_{m}\rrbracket$, and so the chaining argument proceeds by considering a dyadic mesh of points within each interval $I_{m}(j)$, enumerated as follows. For $\ell,m\in\mathbb{N}$, define
@@
so that $|\widetilde{\mathcal{P}}_{\ell,m}|=T_{m}2^{\ell}$. Furthermore, let $\rho_{\ell,m}(t)$ denote the unique element $r\in\widetilde{\mathcal{P}}_{\ell,m}$ minimizing distance to $t$ and satisfying $\pi_{m}(t)=\pi_{m}(r)$ (so that $t$ and $\rho_{\ell,m}(t)$ lie in the same interval $I_{m}(j)$); in particular, $\rho_{0,m}(t)=\pi_{m}(t)$ for each $t\in[0,1]$. Since $Y_{m}(t)$ is continuous on $0,1]\setminus\mathcal{P}_{m}$ and equal to 0 on $\mathcal{P}_{m}$, we have the following for all $t\in[0,1]$:
@@
We now record the key estimate. Consider $\ell,m\in\mathbb{N}$ and $s,t\in\widetilde{\mathcal{P}}_{\ell,m}$ that are consecutive, so that $|s-t|=2^{-\ell}|I_{m}(j)|$ for some $j\leq T_{m}$ (and therefore $\pi_{m}(t)=\pi_{m}(s)$). Then, using the Lipschitz property of cosine/sine as well as Eq.˜2.1, we find
$\mathbb{P}\bigg(|Y_{m}(t)-Y_{m}(s)|>2\sqrt{2}\pi m^{-2}\ell^{-2}\bigg)\\ =\mathbb{P}\bigg(\Big|\sum_{k=2^{m-1}}^{2^{m}-1}X_{k}(t)-X_{k}(s)\Big|>2\sqrt{2}\pi m^{-2}\ell^{-2}\bigg)\leq Ce^{-c\frac{f(m)2^{\ell}}{m^{2}\ell^{2}}},$
where we have used the exponentially-decaying right and left tails of the sum given to us by [36, Chapter III, Theorem 16] (in particular, Eq. (4.3) and (4.6) there). [^2] Note that [36, Chapter III, Lemma 5] allows us to apply Theorem 16 of that chapter, as condition (4.1) is satisfied. For each $m\in\mathbb{N}$, a union bound over $\ell\in\mathbb{N}$ and all possible pairs of $(\rho_{\ell-1,m}(t),\rho_{\ell,m}(t))$ then yields
$\mathbb{P}\Big(\exists_{\ell\in\mathbb{N}}\exists_{t\in[0,1]}:|Y_{m}(\rho_{\ell,m}(t))-Y_{m}(\rho_{\ell-1,m}(t))|>2\sqrt{2}\pi m^{-2}\ell^{-2}\Big)\leq C\sum_{\ell\in\mathbb{N}}T_{m}2^{\ell}e^{-c\frac{f(m)2^{\ell}}{m^{2}\ell^{2}}}.$
Combining the last display with Eq.˜2.13 yields
@@
which, recalling the definition of $Y_{m}(t)$ in Eq.˜2.12 as well as the bound on $T_{m}$ from Eq.˜2.1, implies Eq.˜2.11. This concludes the proof of Proposition˜2.3. ∎
@@ 2.2. Coupling with Gaussian log-correlated fields
In addition to Proposition˜2.3, the other key input for the proof of ˜1.1 is Proposition˜2.7, which gives mutual absolute continuity of $\widetilde{\mu}_{\gamma,\widetilde{a}}$ with a chaos measure associated to a discrete hierarchical Gaussian model constructed on the same probability space. Before stating the proposition, we construct a coupling between $(\widetilde{a}_{n}^{(i)})_{i\in\{1,2\},n\in\mathbb{N}}$ and a Gaussian random vector $(\widetilde{g}_{n}^{(i)})_{i\in\{1,2\},n\in\mathbb{N}}$, where $\widetilde{g}_{n}^{(i)}:=(\widetilde{g}_{n}^{(i)}(v))_{v\in\mathcal{N}_{n}}$, such that the two vectors are close on coordinates that are “thick” (Lemma˜2.6), and
@@
Note Eq.˜2.14 implies “generational independence”: $\widetilde{g}_{n}^{(i)}$ is independent of $\widetilde{g}_{m}^{(i)}$ for $n\neq m$. In fact, we construct i.i.d. standard Gaussians $g_{k}^{(i)}$ such that the $\widetilde{g}_{n}^{(i)}$ are given by taking $a_{k}^{(i)}:=g_{k}^{(i)}$ in (2.3). These will be the $g_{k}^{(i)}$ appearing in the statement of ˜1.1. The construction is completed in (2.24). The main tool here is a variant of the Yurinskii coupling from [5], reproduced below.
Lemma 2.4 ([5, Appendix I, Lemma 38]).
@@
is finite, where each $\mathfrak{g}_{k}$ denotes an independent random vector in $\mathbb{R}^{K}$, $\mathfrak{g}_{k}\sim\mathcal{N}(0,\operatorname{Cov}(\xi_{i}))$. Define $S:=\sum_{k=1}^{n}\xi_{k}$ and let $Z\sim\mathcal{N}(0,I_{K})$. Then, for each $\delta>0$, there exists a random vector $T\sim\mathcal{N}(0,\operatorname{Cov}(S))$ such that
@@
Fix $n\in\mathbb{N}$ and $\mathcal{K}\subseteq\mathcal{N}_{n}$. There exists a $\mathcal{N}(0,\operatorname{Cov}[(\widetilde{a}_{n}^{(1)}(v),\widetilde{a}_{n}^{(2)}(v))_{v\in\mathcal{K}}])$-distributed random vector $\mathfrak{g}_{\mathcal{K},n}$ taking values in $\mathbb{R}^{2|\mathcal{K}|}$ such that
@@
$\mathbb{P}\Big(\big\|\mathfrak{g}_{\mathcal{K},n}-(\widetilde{a}_{n}^{(1)}(v),\widetilde{a}_{n}^{(2)}(v))_{v\in\mathcal{K}}\big\|_{\infty}>n^{-2}\Big)\leq\min_{t\geq 0}\Big\{2\mathbb{P}\Big(\|Z\|_{\infty}>t\Big)+Cn^{6}\beta t^{2}\Big\},$ (2.17)
with $Z=(Z_{j})_{j\leq K}$ i.i.d., $\mathcal{N}(0,1)$-distributed and $\beta$ as in Eq.˜2.15. To control the $\xi_{k}$ part of $\beta$, we use that $\|\xi_{k}\|_{2}^{2}\leq K\|\xi_{k}\|_{\infty}^{2}$ and $\|\xi_{k}\|_{\infty}\leq N^{-1/2}\max\{|a_{l_{n}(k)}^{(1)}|,|a_{l_{n}(k)}^{(2)}|\}$ to get
$\sum_{k=1}^{N}\mathbb{E}[\|\xi_{k}\|_{2}^{2}\,\|\xi_{k}\|_{\infty}]\leq\mathbb{E}[\max\{|a_{1}^{(1)}|,|a_{1}^{(2)}|\}]\cdot N^{-1/2}K.$ (2.18)
The $\mathfrak{g}_{k}$ part of $\beta$ is bounded as $CN^{-1/2}K\log(K)^{1/2}$ by [12, Lemma B.3], so that
@@
Furthermore, using Borel-TIS and $\mathbb{E}\left[\|Z\|_{\infty}\right]\leq C\log(K)^{1/2}$, we get for $t\geq C\log(K)^{1/2}$
$\mathbb{P}\big(\|Z\|_{\infty}>t\big)\leq Ce^{-\frac{(t-C\log(K)^{1/2})^{2}}{2}}$
where we recall the constant $C$ can differ from instance to instance. Plugging this into Eq.˜2.17 bounds the left-hand side of Eq.˜2.17 by
$C\min_{t\geq C\log(K)^{1/2}}\Big\{e^{-\frac{(t-C\log(K)^{1/2})^{2}}{2}}+n^{6}t^{2}N^{-\frac{1}{2}}K\log(K)^{\frac{1}{2}}\Big\}.$
Taking $t=C\log(K)^{1/2}+n$ and recalling $N^{-1/2}=2^{-(n-1)/2}$ and $K=2|\mathcal{K}|$ yields Eq.˜2.16. ∎
We now use Lemma 2.5 to recursively construct the Gaussians $(\widetilde{g}_{n}^{(i)})_{i\in\{1,2\},n\in\mathbb{N}}$. As a first step we set $(\widetilde{g}_{1}^{(i)})_{i\in\{1,2\}}:=\mathfrak{g}_{\mathcal{N}_{1},1}$ from Lemma 2.5.
Assume we have constructed $(\widetilde{g}_{k}^{(i)})_{i\in\{1,2\},k\leq n}$. We construct the $(n+1)^{\mathrm{st}}$ generation as follows. First, we only apply the coupling on points satisfying some notion of thickness “up to level nn”, thereby reducing the quantity of random variables we need to couple. Towards this, fix $\delta>0$ satisfying $\gamma-\delta>1$, and define the following sets of time-$n$ thick points:
$\mathcal{T}_{n,\gamma,\widetilde{a}}$ $:=\left\{v\in\mathcal{N}_{n}:\widetilde{S}_{n,\widetilde{a}}(v)\geq\log(2)(\gamma-\delta)n\right\}\,$,
$\mathcal{T}_{n,\gamma,\widetilde{g}}$ $:=\left\{v\in\mathcal{N}_{n}:\widetilde{S}_{n,\widetilde{g}}(v)\geq\log(2)(\gamma-\delta)n\right\}\,$,
@@
and for all $\delta>0$
$\mathbb{P}\Big(\sup_{v\in\mathcal{N}_{n+1}(\mathcal{T}_{n,\gamma}),i\in\{1,2\}}\Big|\widetilde{g}_{n+1}^{(i)}(v)-\widetilde{a}_{n+1}^{(i)}(v)\Big|>n^{-2}\Big)\leq C2^{-\frac{(\gamma-\delta)^{2}-1}{4+\delta}}n\,.$ (2.20)
Proof.
Observe that by construction $\mathcal{N}_{n+1}(\mathcal{T}_{n,\gamma})$ is independent of $(\widetilde{a}_{n+1}^{(i)})_{i\in\{1,2\}}$. As a consequence, we may apply Lemma˜2.5 to construct a Gaussian random vector $(\widetilde{g}_{n+1}^{(1)}(v),\widetilde{g}_{n+1}^{(2)}(v))_{v\in\mathcal{N}_{n+1}(\mathcal{T}_{n,\gamma})}$ with
$\operatorname{Cov}[(\widetilde{g}_{n+1}^{(1)}(v),\widetilde{g}_{n+1}^{(2)}(v))_{v\in\mathcal{N}_{n+1}(\mathcal{T}_{n,\gamma})}]=\operatorname{Cov}[(\widetilde{a}_{n+1}^{(1)}(v),\widetilde{a}_{n+1}^{(2)}(v))_{k\in\mathcal{N}_{n+1}(\mathcal{T}_{n,\gamma})}]$
@@
On the event $|\mathcal{N}_{n+1}(\mathcal{T}_{n,\gamma})|\leq|\mathcal{N}_{n+1}|\,2^{-\frac{1+(\gamma-\delta)^{2}}{4}n}$, the right-hand side of Eq.˜2.21 is smaller than the right-hand side in Eq.˜2.20, where we made use of $|\mathcal{N}_{n+1}|\leq(n+1)^{4}2^{n+2}$. Furthermore, a first moment calculation yields
@@
We combine Eq.˜2.21 and Eq.˜2.22 to get Eq.˜2.20. ∎
Having constructed the Gaussian vector $(\widetilde{g}_{n}^{(i)}(v))_{v\in\mathcal{N}_{n}(\mathcal{T}_{n-1,\gamma})}$ for each $n\in\mathbb{N}$ and $i\in\{1,2\}$, we now extend this vector to $(\widetilde{g}_{n}^{(i)}(v))_{v\in\mathcal{N}_{n}}$ in such a way that Eq.˜2.14 is satisfied. Using Borel-Cantelli and a first moment bound on the right tail of $|\mathcal{N}_{n}(\mathcal{T}_{n-1,\gamma})|$ there is a random $n_{0}(\omega)$ such that for all $n\geq n_{0}(\omega)$ we can choose $\overline{\mathcal{N}_{n}}$ such that $|\overline{\mathcal{N}_{n}}|=2^{n-1}$ and
@@
First, we describe our coupling for $n\leq n_{0}(\omega)-1$. In this case, we take an independent coupling, i.e., given $n_{0}(\omega)$ we take $(g_{k}^{(1)},g_{k}^{(2)})_{k\leq 2^{n(\omega)-1}-1}$ to be an i.i.d. sequence of $\mathcal{N}(0,1)$ distributed random variables and define $((\widetilde{g}_{n}^{(i)}(v))_{v\in\mathcal{N}_{n}}))_{n\leq n_{0}(\omega)-1}$ analogously to Eq.˜2.3.
Next, we give our coupling for $n\geq n_{0}(\omega)$. In this case, we extend $(\widetilde{g}_{n}^{(i)}(v))_{v\in\mathcal{N}_{n}(\mathcal{T}_{n-1,\gamma})}$ to a Gaussian vector $(\widetilde{g}_{n}^{(i)}(v))_{v\in\overline{\mathcal{N}_{n}}}$ such that
@@
We see from Eq.˜2.3 that Eq.˜2.14 is satisfied. We also note that the $g_{k}^{(i)}$ constructed in Eq.˜2.23 are the ones appearing in the statement of ˜1.1.
As mentioned at the start of the subsection, the relationship between the $g_{k}^{(i)}$ and the $\widetilde{g}_{n}^{(i)}$ is the same as that of the $a_{k}^{(i)}$ and the $\widetilde{a}_{n}^{(i)}$, given by Eq.˜2.3. As such, our construction fits in the framework of Section˜2.1. By taking $a_{k}^{(i)}=g_{k}^{(i)}$ in Eq.˜2.3, we may define the field $\widetilde{S}_{k,g}(v)$ as in Eq.˜2.6, as well as the limiting chaos measure $\widetilde{\mu}_{\gamma,\widetilde{g}}$ as in Eq.˜2.8.